Pure copper is produced by the reaction of copper(
Answer #1

40)
mass of Cu = 0.640 kg = 640 g
molar mass of Cu = 63.546 g/mol
number of moles of Cu = mass/molar mass
= 640/63.546
= 10.07 mol

from reaction gives,
moles of Cu2S required = moles of Cu/2
= 10.07/2
= 5.035 mol

molar mass of Cu2S = 159.16 g/mol

mass of Cu2S = number of moles * molar mass
= 5.035 * 159.16
= 801 g
= 0.801 kg
Answer: b

I am allowed to answer only 1 question at a time